OpenStudy (anonymous):

For the function 1) find lim(x-->a-) and lim(x-->a+) whre x=a is a vertical retricemptote and 2) FInd

OpenStudy (anonymous):

asymptote only has one s.

OpenStudy (anonymous):

Otherwise you still need a definition for the function in order to determine the limits at each side of a.

OpenStudy (anonymous):

\[\lim \lim_{x \rightarrow \infty} f(x)\]

OpenStudy (anonymous):

how is f defined?

OpenStudy (anonymous):

and lim lim_{x rightarrow -infty} f(x) and identify a horizontal retricemptote \[f(x)=(\sqrt{81x^4+9x^2}+3x^2)/3x^3-12)\]

OpenStudy (anonymous):

asymptote really only has one s.